# Sections of Lagrangian fibrations on holomorphic symplectic manifolds

### v4

@@ 3.1 Constructing a smooth section over a curve using Néron models
Step 3: Applying a degenerate twistor deformation, we can always assume that $M_{1}$ is algebraic; indeed, the set of algebraic members of the degenerate twistor deformation is dense in the set of Kähler fibers ([V1]). Let $U\subset M_{1}$ be the smooth locus of $\pi$. Locally over an open subset $W_{C}\subset C$, we can choose a holomorphic section of this fibration, because the fibers of $\pi:\;M_{1}\to C$. Then, by 2.4, the smooth locus of $\pi^{-1}(W_{C})$ is locally identified with its Néron model. The Néron model is a group fibration, with general fiber an abelian variety, and with special fiber an abelian complex group, not necessarily compact. The fiberwise group structure obtained this way is continuous by 2.4. The singular locus $E:=M_{1}\backslash U$ has complex codimension $\geqslant 2$, hence by transversality we can always assume that the manifold $S$ does not intersect $E$. Indeed, the singular locus of $\pi{\left|{}_{{\phantom{|}\!\!}_{M_{1}}}\right.}$ has codimension 1 in the set $\pi^{-1}(C\cap\Delta)$, where $\Delta$ is the discriminant divisor of $\pi$. Since the group structure on the Néron model is continuous, the operation defined in Step 2 can be extended to $S$ considered as a multisection of the Néron model projection to $W_{C}$. This defines a smooth local section of $\pi:\;M_{1}\to C$ over $W_{C}$.
Step 4: It remains to show that the sections that we constructed locally over $W_{C}\subset C$ agree and define a section of $\pi:\;M_{1}\to C$. The possible ambiguity arises from the local choice of the zero section, used in the construction of the Néron model. Clearly, it suffices to check that they agree on the smooth locus of $\pi$. However, outside of the singularities, the fibers of $\pi$ are tori, equipped with an action of the corresponding abelian group, and the operation (3.1) is independent of the choice of the zero.
Remark 3.3: This result was conjectured in [KV, Remark 4.2].
@@ 3.2 Constructing a smooth section over a curve (without Néron models)
We give another proof of 3.1 using the idea communicated to us (after the first draft of this paper was published) by János Kollár. This way, we can get rid of the Néron models in 3.1; this also makes it unnecessary to assume that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
Theorem 3.4:  Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in integer homology. Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Proof. Step 1: Let $M_{1}:=\pi^{-1}(C)$. Like in the proof of 3.1, Step 2, we can assume that $M_{1}$ contains a smooth submanifold $Z$ such that $[Z]\cap F=1$, where $F=\pi^{-1}(z)$ for a general point $z\in C$. Let $F_{1}$ be a singular fiber of $\pi$. In the first and second step, we show that $Z$ can be chosen in such a way that for some neighbourhood $U\supset F_{1}$, $Z\cap U$ is a disconnected union of holomorphic subvarieties (possibly with an opposite orientation).
If $F_{1}$ is reduced, this observation is clear: we deform $Z$ in such a way that it intersects $F_{1}$ transversally in its smooth points, and notice that every smooth section of the projection $(z_{1},z_{2},...,z_{n})\mapsto z_{1}$ is homotopic to a holomorphic section near a point $(0,z_{2},...,z_{n})$.
@@
Step 3: Let $Z_{0}$ be the intersection of $Z$ with the set of all smooth fibers of $\pi:\;M_{1}\to C$. These fibers are complex tori, hence we can apply to $Z_{0}$ the averaging procedure defined in 3.1, Step 2, obtaining a section of $\pi:\;M_{1}\to C$ over the set $C_{0}$ of all regular values of this map. Since near the critical values $Z$ was a holomorphic multivalued function, and the averaging procedure defines a meromorphic map $M_{1}\times_{\mathbb{C}}M_{1}\times_{C}...\times_{C}M_{1}\to M_{1}$, we obtain a section $S$ of $\pi:\;\pi^{-1}(C)\to C$ which is meromorphic near all singular fibers. Since the fibers of $\pi$ are compact, $S$ is holomorphic at each critical value of $\pi:\;\pi^{-1}(C)\to C$ by valuative criterion of properness. Therefore this section is smooth on $C$.
@@ 4.1 Barlet spaces: an introduction
Definition 4.1: The Douady space is the moduli of equidimensional dimensional complex subspaces of $M$ (“subspaces” here means that we allow nilpotents in the structure sheaf). The Hilbert-Chow morphism is the forgetful map $F:\;{\cal D}{\>\longrightarrow\>}{\cal B}$ from the Douady space to the Barlet space, with multiplicity of each component of $\varphi(\widetilde{Z})$ given by the rank of the structure sheaf of $Z$ in the general point of this component.
@@ 4.2 Rationally connected varieties
Definition 4.10:  Let $M$ be an irreducible compact complex variety $M$ bimeromorphic to a compact Kähler manifold. Then $M$ is called rationally (chain) connected if for any proper subvariety $E\subset M$, any two points in $M$ can be connected by a sequence of intersecting rational curves which intersect in points of $M\backslash E$.
@@ 4.3 Holography principle with Kähler target space
Remark 4.13: For 4.3, Kählerness of $X$ is sufficient, but not necessary. As seen from the proof, for 4.3 to hold it suffices to establish that all components of the Barlet spaces of curves on $X$ and $M$ are compact. This holds true if $X$ admits a Hermitian form $\omega$ which is pluriclosed (i.e., it satisfies $dd^{c}\omega=0$), or such that $-dd^{c}\omega$ is a positive form ([V3]).
Remark 4.14: It is not difficult to find a counterexample to 4.3 when $X$ is not simply connected and non-Kähler. For instance, let $X=\operatorname{Tw}(Y)$ be the twistor space for a 2-dimensional complex torus $Y$ (see e.g. [V3]). This is an example of a complex manifold with Barlet components of rational curves non-compact (indeed, in [V3] it is shown that all of these components are Stein). The universal cover of $\operatorname{Tw}(Y)$ is the twistor space for ${\mathbb{C}}^{2}={\mathbb{H}}$, which can be identified with the total space of ${\cal O}(1)^{2}$ ([JV]). Since $\operatorname{Tw}(S^{4})$ is ${\mathbb{C}}P^{3}$ and ${\mathbb{C}}^{2}$ is conformally equivalent to $S^{4}$ without a point, the space $\operatorname{Tw}({\mathbb{C}}^{2})$ is biholomorphic to ${\mathbb{C}}P^{3}\backslash{\mathbb{C}}P^{1}$ (see [JV] for another argument leading to this conclusion). Then an open neighbourhood of ${\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{3}$ can be embedded into $\operatorname{Tw}(Y)$. However, if this embedding can be extended to the whole ${\mathbb{C}}P^{3}$, the space $\operatorname{Tw}(Y)$ would be Moishezon, being an image of a proper holomorphic map from a Moishezon space (see e.g. [F]). This is impossible, because the algebraic dimension of any twistor space of a compact hyperkähler manifold is one ([V2]).
@@
When $X$ is Moishezon, 4.3 is implied by the following theorem (see also [V2, Theorem 3.4]; if we deal with formal schemes instead of complex manifolds, this result was established in [Har] and in [HM, Hir]). The following result requires the manifold $M$ to be homogeneous and rational; for non-homogeneous case, see [V2, Theorem 3.4].
@@
Proof of 4.3: The main idea of the proof of 4.3 is the following observation. Consider an irreducible complex subvariety $B_{0}$ in the Barlet space of curves on $X$ containing $n\varphi(S^{\prime})$, where $n$ is a positive integer (representing multiplicity) and $S^{\prime}\subset U_{S}$ a general deformation of $S$ in $U_{S}$. The corresponding Barlet locus $L_{0}$ is defined as the union of all curves $C\in B_{0}$. Assume that $L_{0}$ is rationally connected. Then $L_{0}$ is Moishezon ([C1]) and its normalization is simply connected ([Deb, Corollary 2.32]; see also [C2, KMM]), hence 4.3 can be applied, showing that the map $\varphi:\;U_{S}{\>\longrightarrow\>}L_{0}$ can be extended to a meromorphic map $M\dashrightarrow L_{0}\subset X$.
Step 1: We take $n=3$. Let $B_{0}$ be the set of all deformations of $3\varphi(S)$ (that is, $\varphi(S)$ with multiplicity 3) which intersect $\varphi(S)$ and intersect $\varphi(U_{S})$ in an open set. The first condition is closed. To get a hold on the second condition, denote by $s$ the intersection of a curve $S_{1}\in B_{0}$ and $\varphi(S)$. Clearly, $S_{1}\cap\varphi(U_{S})$ is open if and only if $S_{1}$ is tangent of infinite order [^1] That is, its $k$-jet in $s$ belongs to the set of $k$-jets of curves in $U_{S}$, for all $k$. to $U_{S}$ in $s$. This condition is closed, because it is obtained by taking an intersection of closed subvarieties enumerated by the order of tangency $k=1,2,3,\dots$. Therefore, $B_{0}$ is closed in the Barlet component $B$ of all deformations of $3\varphi(S)$. From Bishop’s compactness theorem it follows that $B_{0}$ is compact. From Remmert’s proper map theorem, we obtain that the corresponding Barlet locus $L_{0}\subset X$ is complex analytic. To finish the proof, it remains to show that $L_{0}$ is rationally connected. We will prove that two generic points of $L_{0}$ are connected by a smooth curve which belongs to $B_{0}$; by 4.2, this implies the rational connectedness of $L_{0}$.
@@
A union of two deformations $S_{1}\ni a,S_{2}\ni b$ of $S$ which intersect can be deformed to a smooth deformation $S_{3}$ of $2S$ by the standard technique ([Ko1]). Moreover, we can perform this deformational smoothing in such a way that $S_{3}$ contains $a$ and $b$. This is the idea used to prove 4.2.
@@
Step 3: Let $L_{0}\subset X$ be the subvariety constructed in Step 1. By construction, $\varphi(U_{S})$ is an open neighbourhood of $S\subset L_{0}$. Therefore, to prove that $L_{0}$ is rationally connected, it would suffice to find a smooth rational curve in $L_{0}$ connecting two general points in $\varphi(U_{S})$. This was done in Step 2. As shown in Step 1, this implies that $L_{0}$ is rationally connected, and finishes the proof of 4.3.

### v5

@@ 4.1 Constructing a smooth section over a curve using Néron models
Step 3: Applying a degenerate twistor deformation, we can always assume that $M_{1}$ is algebraic; indeed, the set of algebraic members of the degenerate twistor deformation is dense in the set of Kähler fibers ([V1]). Let $M_{2}\subset M_{1}$ be the smooth locus of $\pi$. Locally over an open subset $W_{C}\subset C$, we can choose a holomorphic section of this fibration, because the fibers of $\pi:\;M_{1}\to C$. Then, by 3.1, the smooth locus $M_{2}$ of $\pi^{-1}(W_{C})$ is identified with its torsor Néron model. The Néron model is a torsor of a smooth abelian group fibration, with general fiber a compact complex torus, and with special fiber a torsor of an abelian complex group, not necessarily compact. The fiberwise torsor structure obtained this way is holomorphic. The singular locus $E:=M_{1}\backslash M_{2}$ has complex codimension $\geqslant 2$, hence by transversality we can always assume that the manifold $S$ does not intersect $E$. Indeed, the singular locus of $\pi{\left|{}_{{\phantom{|}\!\!}_{M_{1}}}\right.}$ has codimension 1 in the set $\pi^{-1}(C\cap\Delta)$, where $\Delta$ is the discriminant divisor of $\pi$. Since the torsor structure on $M_{2}$ obtained in 3.1 is continuous, the operation defined in Step 2 can be extended to $S$ considered as a multisection of the projection to $W_{C}$. This defines a smooth local section of $\pi:\;M_{1}\to C$ over $W_{C}$.
Remark 4.3: This result was conjectured in [KV1, Remark 4.2].
4.2 Homology classes of sections of abelian fibrations over a line extended to the ambient manifold
In 4.1 (i) we proved that there exists a smooth section of a Lagrangian fibration restricted to a line ${\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{n}$. This section is obtained by taking a multisection of $M_{1}{\>\longrightarrow\>}{\mathbb{C}}P^{1}$ and applying the averaging procedure as explaned in Subsection 4.1, Step 2.
After an earlier version of this paper was finished, J. Kollár contacted us remarking that this procedure might change the homology class of the section (4.2). This concern is well founded! In the paper [KS], J. Kollár and G. Saccá studied universal compactified Jacobian $\pi:\;M\to{\mathbb{C}}P^{n}$ of a family of curves on a projective surface. This universal compactified Jacobian is a projective variety fibered over a line system of divisors in $S$ (which are curves) with a general fiber over $C\subset S$ identified with the Jacobian $\operatorname{Jac}(C)$. A special case of such an abelian fibration is the classical Beauville-Mukai system on a K3 surface (Subsection 1.5). Fix a general line $L={\mathbb{C}}P^{1}\subset{\mathbb{C}}P^{n}$, and let $M_{1}:=\pi^{-1}(L)$. J. Kollár and G. Saccá study the restriction map $r:\;H^{2g}(M,{\mathbb{Z}})\to H^{2g}(M_{1},{\mathbb{Z}})$, and show that is not surjective. Indeed, for any holomorphic section $s$ of $M_{1}\to L$, such that its fundamental class belongs $\operatorname{im}(r)$, $s$ is necessarily a zero section of $M_{1}\to L$.
In an email letter, J. Kollár gave the following remarkable and unexpected example.
Example 4.4: Let $\pi:\;X\to{\mathbb{C}}P^{1}$ be an elliptic fibration on a K3 surface. Assume that a special fiber of $\pi$ is a collection of at least 3 rational (-2)-curves, and let $\eta]\in H_{2}(X,{\mathbb{Z}})$ be a homology class of a section. Denote by $Z_{1}],...,[Z_{k}]\in H_{2}(X,{\mathbb{Z}})$ the homology classes of the distinct exceptional (-2)-components $Z_{i}$ of a special fiber $F$, and let $n_{1},...,n_{k}$ be a collection of integers. Then $A:=\eta+\sum_{i}n_{i}[Z_{i}]$ is a homology class whose intersection with a general fiber is 1. Using Thom’s theorem as in Subsection 4.1, Step 2, we can represent $A$ by a smooth 2-dimensional submanifold of $X$. The averaging procedure, applied to this situation, produces a section $S$ of the corresponding Néron model, which intersects only one of the irreducible components of $F$, and $A\cap[Z_{i}]=n_{i}[Z_{i}]^{2}+\eta\cap[Z_{i}]$, while $S$, being a section, might intersect only one of $Z_{i}$, with multiplicity 1.
Remark 4.5: It is not clear how to express the fundamental class of $S$ even if we know $A$ and the numbers $A\cap[Z_{i}]$; this seems to be an interesting question.
Proof of 4.1 (ii), Step 1: Let $\alpha\in H_{2n}(M,{\mathbb{Z}})$ be an integer homology class which intersects a general fiber of $\pi:M\to CP^{n}$ in 1. We represent this homology class by a simplicial chain $A$ with smooth simplices transversal to $\pi$. Consider the set $\operatorname{Discr}_{1}\subset{\mathbb{C}}P^{n}$ defined in 3.2, and let $X_{\varepsilon}\subset{\mathbb{C}}P^{n}$ be the complement of its closed $\varepsilon$-neighbourhood. We fix a choice of $C$ and $\varepsilon>0$ such that $C\subset X_{\varepsilon}$.
Recall that Borel-Moore homology is the homology theory dual to cohomology with compact support. The homology classes in Borel-Moore homology are represented by locally finite simplicial chains. Clearly, $A\cap\pi^{-1}(X_{\varepsilon})$ is homologous to a simplicial chain in Borel-Moore homology, and for an appropriate simplicial decomposition of $X_{\varepsilon}$, we can assume that all simplices of $A_{\varepsilon}$ are mapped to the corresponding simplices of $X_{\varepsilon}$.
Let ${\mathfrak{X}}_{\varepsilon}$ be a simplicial decomposition of $X_{\varepsilon}$ such that each simplex of $A_{\varepsilon}$ is projected to a certain simplex of ${\mathfrak{X}}_{\varepsilon}$.
Step 2: By 3.1, the smooth locus of $\pi^{-1}(X_{\varepsilon})$ is an abelian torsor over $X_{\varepsilon}$. Since all simplices of $A_{\varepsilon}$ are transversal to the fibers of $\pi$, and $A_{\varepsilon}\cap\pi^{-1}(pt)=1$ in homology, [^1] Here $A_{\varepsilon}\cap\pi^{-1}(pt)=1$ refers to the cap-product in homology. the intersection of $A_{\varepsilon}$ and $\pi^{-1}(x)$ is 1, for a general $x$, if counted with multiplicities. If we apply the averaging procedure (3.1) using the torsor structure, we obtain that any open simplex $\Delta$ of ${\mathfrak{X}}_{\varepsilon}$ corresponds to an open simplex $s(\Delta)$ obtained by averaging of all simplices of $A$ mapped to $\Delta$. It remains to show that this gives a simplicial (hence, continuous) section of $\pi:\;\pi^{-1}(X_{\varepsilon})\to X_{\varepsilon}$.
Step 3: Let $M_{\varepsilon}:=\pi^{-1}(X_{\varepsilon})$, and $N_{\varepsilon}\subset M_{\varepsilon}$ be the smooth locus. The averaging map
$\underbrace{N_{\varepsilon}\times_{X_{\varepsilon}}...\times_{X_{\varepsilon}}N_{\varepsilon}}_{\text{$n+1$ times}}\times_{X_{\varepsilon}}\underbrace{N_{\varepsilon}\times_{X_{\varepsilon}}...\times_{X_{\varepsilon}}N_{\varepsilon}}_{\text{$n$ times}}\to N_{\varepsilon}$
is holomorphic, hence Lipschitz in any compact neighboorhood of $A$. All simplices $\Delta$ of the chain $A_{\varepsilon}$ can be chosen smooth, hence Lipschitz. Therefore, we can continuously extend each open simplex $s(\Delta)$ to the boundary, obtaining a simplicial chain which gives a simplicial section $\Sigma$ of $\pi$. Intersecting this simplicial section with $\pi^{-1}(C)$ where $C$ is a line, we obtain a homology class satisfying the assumptions of 4.1 (ii), as explained in Step 4 below.
Step 4: The section $\Sigma$ is obtained by taking the homology class $A$ in $H_{2n}(M,{\mathbb{Z}})$, restricting it to the Borel-Moore homology class in $\pi^{-1}(X_{\varepsilon})$, taking the average with respect to the torsor bundle structure (which does not change the Borel-Moore homology class), and restricting the result to the preimage of the curve $C\subset{\mathbb{C}}P^{n}$ by taking the cap-product with $\pi^{-1}(C)$. By construction, this is the same as taking a homology class $A\in H_{2n}(M,{\mathbb{Z}})$, and restricting it to $\pi^{-1}(C)$ by taking the cap-product.
@@ 4.3 Constructing a smooth section over a curve (without Néron models)
We give another proof of 4.1 (i) using the idea communicated to us (after the first draft of this paper was published) by János Kollár. This way, we can get rid of the Néron models in 4.1 (i); this also makes it unnecessary to assume that $\pi:\;\pi^{-1}(C)\to C$ has local sections.
Theorem 4.6: Let $\pi:\;M\to X$ be a hyperkähler manifold equipped with an Lagrangian fibration, with $X$ smooth (and hence, by Hwang’s theorem, isomorphic to ${\mathbb{C}}P^{n}$), and $C\subset X$ a general line in $X={\mathbb{C}}P^{n}$. Assume that the general fiber of the Lagrangian projection $\pi$ is primitive (that is, not divisible) in the integral homology. Then there exists a smooth section $\sigma:\;C{\>\longrightarrow\>}M$.
Proof. Step 1: Let $M_{1}:=\pi^{-1}(C)$. Like in the proof of 4.1, Step 2, we can assume that $M_{1}$ contains a smooth submanifold $Z$ such that $Z]\cap F=1$, where $F=\pi^{-1}(z)$ for a general point $z\in C$. Let $F_{1}$ be a singular fiber of $\pi$. In the first and second step, we show that $Z$ can be chosen in such a way that for some neighbourhood $U\supset F_{1}$, $Z\cap U$ is a disconnected union of holomorphic subvarieties (possibly with an opposite orientation).
If $F_{1}$ is reduced, this observation is clear: we deform $Z$ in such a way that it intersects $F_{1}$ transversally in its smooth points, and notice that every smooth section of the projection $(z_{1},z_{2},...,z_{n})\mapsto z_{1}$ is homotopic to a holomorphic section near the point $(0,z_{2},...,z_{n})$.
@@
Step 3: Let $Z_{0}$ be the intersection of $Z$ with the set of all smooth fibers of $\pi:\;M_{1}\to C$. These fibers are complex tori, hence we can apply to $Z_{0}$ the averaging procedure defined in 3.1, obtaining a section of $\pi:\;M_{1}\to C$ over the set $C_{0}$ of all regular values of this map. Since near the critical values $Z$ was a holomorphic multivalued function, and the averaging procedure defines a meromorphic map $M_{1}\times_{\mathbb{C}}M_{1}\times_{C}...\times_{C}M_{1}\to M_{1}$, we obtain a section $S$ of $\pi:\;\pi^{-1}(C)\to C$ which is meromorphic near all singular fibers. Since the fibers of $\pi$ are compact, $S$ is holomorphic at each critical value of $\pi:\;\pi^{-1}(C)\to C$ by the valuative criterion of properness. Therefore, this section is smooth on $C$.
@@ 5.1 Barlet spaces: an introduction
[…]
@@ 5.2 Rationally connected varieties
[…]
@@ 5.3 Holography principle with Kähler target space
Remark 5.13: For 5.3, Kählerness of $X$ is sufficient, but not necessary. As seen from the proof, for 5.3 to hold it suffices to establish that all components of the Barlet spaces of curves on $X$ and $M$ are compact. This holds true if $X$ admits a Hermitian form $\omega$ which is pluriclosed (i.e., it satisfies $dd^{c}\omega=0$), or such that $dd^{c}\omega$ is a positive form ([V3]).
[…]
@@
When $X$ is Moishezon, 5.3 is implied by the following theorem (see also [V2, Theorem 3.4]; if we deal with formal schemes instead of complex manifolds, this result was established in [Har] and in [HM, Hir]). The following result requires the manifold $M$ to be homogeneous and rational; for non-homogeneous case, see [V2, Theorem 3.4].
@@
Proof of 5.3: The main idea of the proof of 5.3 is the following observation. Consider an irreducible complex subvariety $B_{0}$ in the Barlet space of curves on $X$ containing $n\varphi(S^{\prime})$, where $n$ is a positive integer (representing multiplicity) and $S^{\prime}\subset U_{S}$ a general deformation of $S$ in $U_{S}$. The corresponding Barlet locus $L_{0}$ is defined as the union of all curves $C\in B_{0}$. Assume that $L_{0}$ is rationally connected. Then $L_{0}$ is Moishezon ([C1]) and its normalization is simply connected ([Deb, Corollary 2.32]; see also [C2, KMM]), hence 5.3 can be applied, showing that the map $\varphi:\;U_{S}{\>\longrightarrow\>}L_{0}$ can be extended to a meromorphic map $M\dashrightarrow L_{0}\subset X$.
Step 1: We take $n=3$. Let $B_{0}$ be the set of all deformations of $3\varphi(S)$ (that is, $\varphi(S)$ with multiplicity 3) which intersect $\varphi(S)$ and intersect $\varphi(U_{S})$ in an open set. The first condition is closed. To get a hold on the second condition, denote by $s$ the intersection of a curve $S_{1}\in B_{0}$ and $\varphi(S)$. Clearly, $S_{1}\cap\varphi(U_{S})$ is open if and only if $S_{1}$ is tangent of infinite order [^1] That is, its $k$-jet in $s$ belongs to the set of $k$-jets of curves in $U_{S}$, for all $k$. to $U_{S}$ in $s$. This condition is closed, because it is obtained by taking an intersection of closed subvarieties enumerated by the order of tangency $k=1,2,3,\dots$. Therefore, $B_{0}$ is closed in the Barlet component $B$ of all deformations of $3\varphi(S)$. From Bishop’s compactness theorem it follows that $B_{0}$ is compact. From Remmert’s proper map theorem, we obtain that the corresponding Barlet locus $L_{0}\subset X$ is complex analytic. To finish the proof, it remains to show that $L_{0}$ is rationally connected. We will prove that two generic points of $L_{0}$ are connected by a smooth curve which belongs to $B_{0}$; by 5.2, this implies the rational connectedness of $L_{0}$.
@@
A union of two deformations $S_{1}\ni a,S_{2}\ni b$ of $S$ which intersect can be deformed to a smooth deformation $S_{3}$ of $2S$ by the standard technique ([Ko1]). Moreover, we can perform this deformational smoothing in such a way that $S_{3}$ contains $a$ and $b$. This is the idea used to prove 5.2.
@@
Step 3: Let $L_{0}\subset X$ be the subvariety constructed in Step 1. By construction, $\varphi(U_{S})$ is an open neighbourhood of $S\subset L_{0}$. Therefore, to prove that $L_{0}$ is rationally connected, it would suffice to find a smooth rational curve in $L_{0}$ connecting two general points in $\varphi(U_{S})$. This was done in Step 2. As shown in Step 1, this implies that $L_{0}$ is rationally connected, and finishes the proof of 5.3.